# VCR: Visual Caption Restoration

### v3

@@ Abstract
We introduce Visual Caption Restoration (VCR), a novel vision-language task that challenges models to accurately restore partially obscured texts using pixel-level hints within images. This task stems from the observation that text embedded in images intrinsically differs from common visual elements and natural language due to the need to align the modalities of vision, text, and text embedded in images. While numerous works have integrated text embedded in images into visual question-answering tasks, approaches to these tasks generally rely on optical character recognition or masked language modeling, thus reducing the task to mainly text-based processing. However, text-based processing becomes ineffective in VCR as accurate text restoration depends on the combined information from provided images, context, and subtle cues from the tiny exposed areas of masked texts. We develop a pipeline to generate synthetic images for the VCR task using image-caption pairs, with adjustable caption visibility to control the task difficulty. With this pipeline, we construct a dataset for VCR called VCR-wiki using images with captions from Wikipedia, comprising $2.11M$ English and $346K$ Chinese entities in both easy and hard configurations. Our results reveal that current vision language models significantly lag behind human performance in the VCR task, and merely fine-tuning the models on our dataset does not lead to notable improvements. We release VCR-wiki and the data construction code to facilitate future research.
@@ 1 Introduction
Recent advances in large language models, such as ChatGPT  (Ouyang et al., 2022; OpenAI et al., 2023) and Llama (Touvron et al., 2023), have spurred significant interest and progress in the field of vision-language models. With models like GPT-4V (OpenAI et al., 2023) and LLaVA (Liu et al., 2023a; 2024a; b) blending textual and visual information, the intersection of computer vision and natural language processing has become a vibrant research frontier. These integrated models aim to leverage the potential of vision and language modalities to understand and interpret multimedia content more effectively.
@@
The Visual Question Answering (VQA) task (Antol et al., 2015; Wang et al., 2018; Mishra et al., 2019b; Singh et al., 2019) has been a popular benchmark in assessing how well models align and interpret visual and linguistic information. Traditional VQA approaches, however, predominantly focus on direct queries about visible elements in images and do not address the nuanced relationship between textual content embedded within the image and the overall image context. This gap underscores the limited capabilities of current models in processing integrated visual-textual data, particularly when the textual component, which plays a critical role, is partially obscured or altered.
@@
To develop the VCR task, in this work, we introduce a pipeline for generating synthetic images that allows for manipulation of the visibility of the textual components of the image. This not only enhances the challenge posed by the task but also provides a scalable way to adjust task difficulty. The resulting dataset, VCR-wiki, comprises 2.11M English data and 346K Chinese data sourced from Wikipedia, featuring captions in both languages across ‘easy’ and ‘hard’ difficulty levels. Our evaluations indicate that existing vision-language models significantly underperform compared to human benchmarks, underscoring the need for novel model architectures and training paradigms specifically geared towards this complex intermodal alignment.
By releasing VCR-wiki and the accompanying dataset construction code, we aim to stimulate further research in this area, encouraging the development of models that can more adeptly navigate the nuanced landscape of the restoration of text embedded in images. This effort aligns with the broader goal of advancing vision-language models to achieve a deeper, more intuitive understanding of multimedia content, bridging the gap between human and machine perception. The code is available at https://github.com/tianyu-z/VCR.
@@
    Introduce the VCR task to challenge vision-language models to restore occluded texts in images.
@@
    Create and release VCR-wiki, a dataset with multilingual captions from Wikipedia images, designed to benchmark vision-language models (VLMs) on text restoration tasks.
@@
    Conduct empirical evaluations that show significant gaps between current models and human performance on the VCR task. This highlights the effectiveness of VCR for assessing advancements in VLMs, and underscores the necessity for innovative model architectures and training techniques.
@@ 2 VCR Task Description
In this section, we compare the VCR task with other existing tasks and aim to answer the following questions:
@@
For better clarity, we define text embedded in image ($TEI$) as text incorporated within the image. The term visual image (VI) pertains to the portion of the image that excludes the text embedded in the image. The string text (ST) is not part of the image itself but is associated with it as a distinct textual element. It is usually the question prompt in the form of natural language, for example, ‘What are the covered texts in the image? Please only guess the covered texts without outputting an explanation.’. Consequently, an element of a VCR task can be expressed as $(ST,(VI,TEI))$, where $ST$ is represented as a string and both $VI$ and $TEI$ are presented in image form. This notation does not imply that $VI$ and $TEI$ can be physically separated into two distinct image components. Instead, this definition is adopted merely to facilitate a clearer explanation of the concepts involved. Please refer to Figure 3 for an illustration of $VI$, $TEI$, and $ST$.
A1 Existing tasks that are similar to VCR are the tasks of VQA and OCR. VQA takes as input images and a natural language question and generates a free-form response. As the ground-truth response is not unique, evaluating VQA poses a major challenge. In contrast to VQA, OCR is a task where the ground-truth responses are unique: OCR takes as input complete characters in image form and outputs a string representing the characters in the image, without considering the image context. Models pretrained with OCR are able to retrieve texts embedded in the input image, even if they are incomplete or vague. However, as the vagueness or occlusion of the textual components of the image increases, retrieving the original text without considering the remaining nontextual image context becomes harder, and OCR is no longer a good approach. VCR bridges the gap between OCR and VQA: it reconstructs the unique text found in the image while also considering the visual context of the rest of the image.
Figure 3 is an example VCR task in hard mode, and Figure 1 shows an example VCR task in the easy mode. Although humans can still fill the blanks easily in the hard mode, it is nearly impossible for models with only OCR capabilities to recover the covered texts without using the context. This is because the pixel-level hints of single characters no longer correspond to a unique solution.
@@
The first aspect of importance stems from discoveries in neuroscience about human cognitive abilities to recognize partially occluded objects (Fyall et al., 2017; Li et al., 2023a). Although existing models can recognize objects and texts in images, they often struggle with the complexity of occluded objects due to significant information loss in the images. In contrast, humans excel at filling in missing information using a combination of low-level visual processing and high-level cognitive functions, such as those managed by the prefrontal cortex. This cortical area is known to handle complex cognitive processes such as decision making and memory retention, which are essential for integrating fragmented visual input into coherent objects. We believe that the occlusion restoration task serves as a probe that can effectively distinguish low-level recognition and high-level cognition involving reasoning. In addition, understanding these neural mechanisms can inspire new algorithms capable of mimicking human-like perception and interpretation in dynamic, real-world conditions where occlusion is common.
@@
The second aspect underscores the unique challenge presented by the VCR task, distinguishing it significantly from existing benchmarks, such as traditional VQA or the occluded object restoration task. By occluding texts instead of common visual objects, VCR targets the models’ text-image alignment capability, which is one of the major challenges for vision-language models. VCR mandates that models align textual and visual information in a manner that replicates human-like understanding involving the utilization of both textual and visual clues. This task requires a deep integration of visual ($VI$), embedded textual ($TEI$), and contextual interpretation across modalities, pushing beyond simple text extraction as performed in OCR tasks. In OCR, the focus is primarily on recognizing visible characters, often without the need to understand their contextual relevance within the image narrative. In contrast, VCR introduces complexity by requiring the model to use available partial texts and the visual context collaboratively to reconstruct the obscured content accurately. This not only tests the model’s ability to process $TEI$-$VI$ modalities effectively but also challenges it to maintain internal consistency, akin to human cognitive processes where context and visual clues guide understanding and response. Besides, the difficulty of the task can be adjusted by altering the extent of text occlusion, offering a scalable and flexible framework for systematically enhancing model capabilities in text-visual alignment and semantic comprehension. This rigorous testing ground will help evolve vision-language models to better grasp the nuanced interplay between text and imagery. We show an example of how humans would solve the VCR task in ‘hard’ difficulty in Figure 2.
@@ 3 Dataset Creation
Within $TEI$, we use spaCy to randomly select several 5-grams in the caption for masking. To ensure the restoration process is doable by a human without too much professional domain knowledge, the 5-grams do not contain numbers, person names, religious or political groups, facilities, organizations, locations, dates, and time labeled by spaCy, and the total masked token does not exceed 50% of the tokens in the caption. We exclude all instances that do not have a single eligible 5-gram. The selected 5-grams are partially obscured by a white rectangle that reveals only the upper and lower parts of the text, with the proportion of coverage varying to adjust task difficulty. Furthermore, to assess the impact of $VI$ on model performance, we create an ablation for each image, maintaining the resolution of the $VI$ +$TEI$ image, but retaining only the $TEI$ part in the center of the image.
The VCR task involves a predefined question that prompts the model to produce the obscured text in the image. The ground truth answer corresponds to the caption displayed in the uncovered portion of the stacked image. Due to the extensive availability of vision-language models and a significant user base in both English and Chinese, we have chosen to develop the dataset in these two languages. For each language, we meticulously select the covering proportion to create two task variants: (1) an easy version, where the task is easy for native speakers but open-source OCR models almost always fail, and (2) a hard version, where the revealed part consists of only one to two pixels for the majority of letters or characters, yet the restoration task remains feasible for native speakers of the language.
We release the dataset under the CC BY-SA 4.0 license. All configurations of VCR-wiki are publicly available at https://huggingface.co/collections/vcr-org/vcr-visual-caption-recognition-6661393b1761e2aff7b967b9.
@@ 3.1 Dataset Format and Statistics
Table 1: Basic statistics of the dataset. Note that the Easy and Hard configurations for each language share the same statistics. We report the mean, standard deviation, and the 5th and 95th percentile ($\eta_{.5}$ and $\eta_{.95}$) for the stacked image height and the number of obscured text spans. Unit is in pixels.
# Train # Val # Test $VI$ +$TEI$ Image Height # Obscured Text Spans
Mean SD $\eta_{.5}$ $\eta_{.95}$ Mean SD $\eta_{.5}$ $\eta_{.95}$
English 2095733 5000 5000 375.52 106.01 253 564 1.62 0.63 1 3
Chinese 336448 5000 5000 360.44 102.76 239 562 2.06 0.94 1 4
The VCR dataset comprises four configurations: English Easy, English Hard, Chinese Easy and Chinese Hard. Each configuration can be further divided into training, validation, and test splits. The validation and test splits contain 5,000 entities each. The training set for English configurations and Chinese configurations contains $2,095,733$ and $336,448$ instances, respectively, which can be used for model continuous pretraining. We include more detailed statistics of the dataset in Table 1.
@@ 4 Experiments
In this section, we report the experiment results of existing state-of-the-art vision-language models on our proposed VCR tasks. The fine-tuning and evaluation of open-source models are conducted on a mix of NVIDIA A100 80G and L40S 48G GPUs in an internal cluster.
@@ 4.1 Models
Closed-source Models.
We evaluate several most advanced proprietary models with their provided APIs. The evaluated models include GPT-4o (gpt-4o-2024-0513), GPT-4 Turbo (gpt-4-turbo-2024-04-09), GPT-4V (gpt-4-1106-vision-preview) (Ouyang et al., 2022; OpenAI et al., 2023), Claude 3 Opus (claude-3-opus-20240229), Claude 3.5 Sonnet (claude-3-5-sonnet-20240620) (Anthropic, 2024), Gemini 1.5 pro (gemini-1.5-pro-001) (Team et al., 2024a), Reka Core (reka-core-20240501) (Team et al., 2024b), and Qwen-VL-Max (tested in May 2024) (Bai et al., 2023).
Open-source Models.
We evaluate open-source models with the best performance on the OpenVLM Leaderboard [^2] We selected the highest-performing open-source models with fewer than 40 billion parameters from the OpenVLM Leaderboard as of May 23, 2024 and their later versions. Details are available at https://huggingface.co/spaces/opencompass/open_vlm_leaderboard. and state-of-the-art Chinese VLM models. The evaluated models include Cambrian-1 (Tong et al., 2024), CogVLM2-19B (Hong et al., 2024), DeepSeek-VL-7B-Chat (Lu et al., 2024), DocOwl-1.5-Omni (Hu et al., 2024a), Idefics3-8B (Laurençon et al., 2024), InternLM-XComposer2-VL-7B (Dong et al., 2024a), InternLM-XComposer2.5-VL (Zhang et al., 2024), InternLM-XComposer2-VL-4K (Dong et al., 2024b), Llama-3.2 (Dubey et al., 2024), MiniCPM-V2.5 (Hu et al., 2024b), Monkey (Liu et al., 2024b; Li et al., 2023b), Pixtral (Mistral, 2024), Qwen-VL-Chat (Bai et al., 2023), Qwen2-VL (Wang et al., 2024), and Yi-VL (01.AI et al., 2024). Out of these models, Cambrian-1, Idefics3 and Llama-3.2 are English-only models, and CogVLM2-Llama3-19B-Chat has its Chinese variant, CogVLM2-Llama3-19B-Chinese-Chat. Please refer to Table 6 for the model specifications.
Finetuned Models.
@@
More specifically, we fine-tune CogVLM2-Llama3-19B-Chat, MiniCPM-Llama3-V2.5, and Qwen2-VL-7B-Instruct in the English Hard configuration, and CogVLM2-Llama3-19B-Chinese-Chat, MiniCPM-Llama3-V2.5, and Qwen2-VL-7B-Instruct on the Chinese Hard configuration. The models are finetuned using LoRA (Hu et al., 2022) with $r=8$ and $\alpha=32$. We adopt the schedule-free AdamW optimizer (Defazio et al., 2024) with a learning rate $42e-$ 4. The effective batch size is 64. Each model is trained on the first 16,000 examples of the training set for 1 epoch. All fine-tuning experiments are performed on a single node with 4 NVIDIA L40S 48G GPUs.
@@ 4.2 Metrics
we provide a heatmap of the correlation matrix in Appendix B.
$\text{VCR}_{\text{ZH, EASY}}$ and $\text{VCR}_{\text{ZH, HARD}}$ were excluded from these processes due to the limited availability of VLMs that support Chinese. From Figure 4, we observe that $\text{VCR}_{\text{EN, EASY}}$ shows a tentative similarity to ChartQA and TextVQA, as all three benchmarks evaluate the ability to extract and reason about text from natural images and documents. However, $\text{VCR}_{\text{EN, EASY}}$ does not exhibit significant similarity to the other benchmarks. On the other hand, $\text{VCR}_{\text{EN, HARD}}$ stands apart from all other benchmarks. We attribute this to the fact that $\text{VCR}_{\text{EN, HARD}}$ emphasizes caption recovery with minimal pixel-level information, a skill not tested by any of the other benchmarks. Therefore, we assert that the VCR series benchmarks assess unique aspects of VLMs that are not covered by any of the other benchmarks in our evaluation.

### v4

# VCR: A Task for Pixel-Level Complex Reasoning in Vision Language Models via Restoring Occluded Text

@@ Abstract
We introduce Visual Caption Restoration (VCR), a novel vision-language task that challenges models to accurately restore partially obscured texts using pixel-level hints within images through complex reasoning. This task stems from the observation that text embedded in images intrinsically differs from common visual elements and text due to the need to align the modalities of vision, text, and text embedded in images. While many works incorporate text into image for visual question answering, they mostly rely on OCR or masked language modeling, reducing the task to text-based processing. However, text-based processing becomes ineffective in VCR as accurate text restoration depends on the combined information from provided images, context, and subtle cues from the tiny, exposed areas of masked texts. We develop a pipeline to generate synthetic images for the VCR task using image-caption pairs, with adjustable caption visibility to control the task difficulty. With this pipeline, we construct VCR-wiki for VCR using Wikipedia images with captions, including $2.11M$ English and $346K$ Chinese training entities, plus $5K$ validation and $5K$ test entities in both languages, each in easy and hard configurations. We also make a hidden test set VCR-Hidden to avoid potential over-fitting on VCR-wiki. Our results reveal that current vision language models significantly lag behind human performance in the VCR task, and merely fine-tuning the models on our dataset does not lead to notable improvements. We release VCR-wiki and the data construction code to facilitate future research.
@@ 1 Introduction
Recent advances in large language models, such as ChatGPT  (OpenAI et al., 2023) and Llama (Touvron et al., 2023), have spurred significant interest and progress in the field of vision-language models (VLMs). With models like GPT-4V (OpenAI et al., 2023) and LLaVA (Liu et al., 2023a; 2024a; b) blending textual and visual information, the intersection of computer vision and natural language processing has become a vibrant research frontier. These integrated models aim to leverage the potential of vision and language modalities to understand and interpret multimedia content more effectively.
@@
The Visual Question Answering (VQA) task (Antol et al., 2015; Wang et al., 2018; Mishra et al., 2019b; Singh et al., 2019) has been a popular benchmark in assessing how well models align and interpret visual and linguistic information. Traditional VQA mainly addresses visible elements, overlooking the nuanced relationship between embedded text and image context. This highlights the limitations of current models in handling integrated visual-textual data, especially when text is obscured or altered.
@@
To develop the VCR task, in this work, we introduce a pipeline for generating synthetic images that allows for manipulation of the visibility of the textual components of the image. This not only enhances the challenge posed by the task but also provides a scalable way to adjust task difficulty. The resulting dataset, VCR-wiki, comprises 2.11M English data and 346K Chinese data sourced from Wikipedia, featuring captions in both languages across ‘easy’ and ‘hard’ difficulty levels. Our evaluations indicate that existing vision-language models significantly underperform compared to human benchmarks, underscoring the need for novel model architectures and training paradigms specifically geared towards this complex intermodal alignment. We also constructed a hidden test set for the VCR task (VCR-Hidden) to avoid potential over-fitting on VCR-wiki.
We release VCR-wiki and its construction code to stimulate further research of developing of models that can more adeptly navigate the nuanced landscape of the restoration of text embedded in images to bridge the gap between human and machine perception. The code and datasets are available at GitHub and Hugging Face.
@@
    Introduce the VCR task to challenge VLMs to restore occluded texts in images.
@@
    Create and release VCR-wiki, a dataset with multilingual captions and construct the hidden test set VCR-Hidden, designed to benchmark VLMs on text restoration tasks.
@@
    Conduct empirical evaluations that show significant gaps between current models and human performance on the VCR task. This highlights the effectiveness of VCR for assessing advancements in VLMs and underscores the necessity for innovative model architectures and training techniques. New models will be actively added to our Github leaderboard.
@@ 2 VCR Task Description
In this section, we compare the VCR task with other existing tasks and answer the following questions:
@@
For better clarity, we define text embedded in image ($TEI$) as text incorporated within the image, visual image (VI) as the non-textual portion of the image, and string text (ST) as the separate textual element associated with the image (typically the question prompt). A VCR task element can thus be expressed as $(ST,(VI,TEI))$, where $ST$ is a string while $VI$ and $TEI$ are presented in image form. We adopt this notation to facilitate explanation, not to imply physical separation of $VI$ and $TEI$ in the image. Please refer to Figure 3 for an illustration of $VI$, $TEI$, and $ST$.
A1 VCR relates to both VQA and OCR tasks. VQA takes images and questions as input to generate free-form responses with non-unique ground-truth, creating evaluation challenges regarding non-unique answers. In contrast to VQA, OCR is a task where the ground-truth responses are unique: OCR takes as input complete characters in image form and outputs a string representing the characters in the image, without considering the image context. VCR bridges these tasks by reconstructing unique text while considering visual context. Figure 2 shows a hard-mode VCR task where humans can fill blanks easily, but models with only OCR capabilities cannot recover covered text without context, as pixel-level character hints no longer yield unique solutions.
@@
First, it connects to fundamental neuroscience findings on human cognitive abilities to recognize partially occluded objects (Fyall et al., 2017; Li et al., 2023a). While existing models struggle with occluded information, humans excel by combining low-level visual processing with high-level cognitive functions in the prefrontal cortex. VCR serves as a critical probe distinguishing between low-level recognition and high-level reasoning cognition—a distinction essential for advancing AI systems toward human-like perception capabilities.
@@
Second, VCR presents a unique challenge substantially different from existing benchmarks by specifically targeting text-image alignment capabilities. Unlike traditional VQA or occluded object restoration tasks that test general visual reasoning, VCR creates a specialized evaluation framework that tests a model’s ability to maintain semantic consistency across multiple modalities simultaneously. It requires deep integration of visual content ($VI$) with partially visible textual elements ($TEI$), necessitates inference capabilities that go well beyond pattern recognition toward genuine comprehension, and demands contextual reasoning that mirrors human cognitive processes when faced with incomplete information.
What makes VCR particularly valuable as a benchmark is its precision in targeting the specific frontier of vision-language integration. By occluding text rather than objects, VCR creates a controlled environment where success requires sophisticated cross-modal reasoning rather than mere recognition or memorization. The ability to adjust difficulty through varying occlusion levels provides researchers with a finely calibrated instrument to measure incremental progress in model capabilities.
For practitioners developing next-generation multimodal systems, VCR offers several distinct advantages: (1) it provides a reliable measure of text-visual alignment capabilities currently lacking in standard benchmarks; (2) it simulates real-world scenarios where text is partially visible, damaged, or obscured; and (3) it offers clear evaluation metrics with unique ground-truth answers, unlike subjective benchmarks that suffer from evaluation ambiguity. Figure 2 demonstrates how humans solve a hard-mode VCR task, highlighting the cognitive processes that advanced AI systems should aim to replicate.
@@ 3 Dataset Creation
We use spaCy to randomly select 5-grams in the caption for masking. To ensure the restoration process is doable by a human without too much domain knowledge, the 5-grams do not contain numbers, person names, religious or political groups, facilities, organizations, locations, dates, and times labeled by spaCy. The total masked token does not exceed 50% of the tokens in the caption. We pick 5-grams for masking as it balances linguistic complexity and task feasibility, capturing meaningful grammatical structures while avoiding dataset reduction or overly simplified tasks observed with longer or shorter spans. We exclude instances that do not have any maskable 5-grams. The selected 5-grams are partially obscured by a white rectangle that reveals only the upper and lower parts of the text, with the proportion of coverage varying according to task difficulty. Furthermore, to assess the impact of $VI$ on model performance, we create an ablation for each image, maintaining the resolution of the $VI$ +$TEI$ image, but retaining only the $TEI$ part in the center of the image.
The VCR task involves a predefined question that prompts the model to produce the obscured text in the image. The ground-truth answer corresponds to the caption displayed in the uncovered portion of the stacked image. Due to the extensive availability of VLMs and a significant user base in both English and Chinese, we have chosen to develop the dataset in these two languages. For each language, we meticulously select the height of the masking rectangle to create two task variants: (1) an easy version, where the task is easy for native speakers but open-source OCR models almost always fail, and (2) a hard version, where the revealed part consists of only one to two pixels for the majority of letters or characters, yet the restoration task remains feasible for native speakers.
To avoid test data leakage, we create a hidden test (VCR-Hidden) for the VCR task which will not be publicized. While it follows our dataset’s general construction principles regarding masked span length and the span selection criteria, it differs in five key aspects: (1) Image-text pairs no longer come from the Wikipedia; (2) $TEI$ is randomly positioned either above or below $VI$; (3) Masked regions are randomly selected from both “easy” and “hard” settings; (4) Multiple popular fonts are used randomly for the $TEI$ component; and (5) The $VI$ +$TEI$ image width varies uniformly between 600 - 1000 pixels. We include results on VCR-Hidden in Table 4.4 and Appendix C.
@@ 3.1 Dataset Format and Statistics
The VCR-wiki dataset comprises four configurations: English Easy, English Hard, Chinese Easy and Chinese Hard. Each configuration can be further divided into training, validation, and test splits. The validation and test splits contain 5,000 entities each. The training set for English configurations and Chinese configurations contains 2,095,733 and 336,448 instances, respectively, which can be used for model continuous pretraining. To avoid test data leakage, we also newly include hidden test sets (VCR-Hidden) for both languages as described in Section 3, which contains 100 entities for each language. We include detailed statistics of the dataset in Table 4 in Appendix A.
@@ 4 Experiments
In this section, we report the experimental results of existing state-of-the-art vision-language models on both VCR-wiki and the VCR-Hidden hidden test.
@@ 4.1 Models
Closed-source and Open-source Models.
In this paper, we report results for several state-of-the-art closed-source and open-source models from the OpenVLM Leaderboard [^2] We selected the highest-performing open-source models with fewer than 40 billion parameters from https://huggingface.co/spaces/opencompass/open_vlm_leaderboard as of May 2024 and their later versions for VCR-wiki., as well as selected state-of-the-art models as of February 2025 for VCR-wiki. For the VCR-Hidden hidden test, we report results of state-of-the-art closed-source and open-source models available as of February 2025. See Appendix B for complete model specifications. We commit to evaluating emerging state-of-the-art VLMs to reflect cutting-edge advancements. Results for later models will be actively updated in the leaderboard hosted on https://github.com/tianyu-z/VCR.
Fine-tuned Models.
@@
More specifically, we fine-tune CogVLM2-Llama3-19B-Chat, MiniCPM-Llama3-V2.5, and Qwen2-VL-7B-Instruct in the English Hard configuration, and CogVLM2-Llama3-19B-Chinese-Chat, MiniCPM-Llama3-V2.5, and Qwen2-VL-7B-Instruct on the Chinese Hard configuration. The models are finetuned using LoRA (Hu et al., 2022) with $r=8$ and $\alpha=32$. We adopt the schedule-free AdamW optimizer (Defazio et al., 2024) with a learning rate $2\mathrm{e}{-4}$. The effective batch size is 64. Each model is trained on the first 16,000 examples of the training set for 1 epoch. All fine-tuning experiments are performed on a single node with 4 NVIDIA L40S 48G GPUs.
@@ 4.2 Metrics
we provide a heatmap of the correlation matrix in Appendix E.
$\text{VCR}_{\text{ZH, EASY}}$ and $\text{VCR}_{\text{ZH, HARD}}$ were excluded from these processes due to the limited availability of VLMs that support Chinese. According to Figure 4, $\text{VCR}_{\text{EN, EASY}}$ shows a tentative similarity to ChartQA and TextVQA, as all three benchmarks evaluate the ability to extract and reason about text from natural images and documents. However, $\text{VCR}_{\text{EN, EASY}}$ does not exhibit significant similarity to the other benchmarks. Meanwhile, $\text{VCR}_{\text{EN, HARD}}$ stands apart from all other benchmarks. We attribute this to the fact that $\text{VCR}_{\text{EN, HARD}}$ emphasizes caption recovery with minimal pixel-level information, a skill not tested by any of the other benchmarks. Therefore, we assert that the VCR series benchmarks assess unique aspects of VLMs that are not covered by any of the other benchmarks in our evaluation.